OpenStudy (anonymous):

Fan & medal x+1/4 = 8/9 Please explain how to do it, I need to lean this...

OpenStudy (anonymous):

\[x + \frac{1}{4} = \frac{8}{9}\]

OpenStudy (anonymous):

first you need to move the 1/4 to the other side of the equation

sammixboo (sammixboo):

To solve for x, you need to make x alone on one side of the equation. start this off my subtracting 1/4 on both sides of the equation, since you are originally adding 1/4 x + 1/4 = 1/4 = 8/9 - 1/4 x = 8/9 - 1/4 Remember that 1/4 - 1/4 = 0, so that is why it is cancled out in the second step

OpenStudy (anonymous):

\[x = \frac{8}{9} - \frac{1}{4}\]

OpenStudy (anonymous):

I think I typed it wrong x+1 is the numerator and 4 is the denminator

OpenStudy (triciaal):